# An adaptive finite element multigrid solver using GPU acceleration

### v1

@@ 5.1 Transport-diffusion equation
Refer to captionRefer to caption
Figure 3: Transport-diffusion equation: Wall-clock times CPU vs. GPU on the finest mesh level with 1 000 00010000001\,000\,0001 000 000 degrees of freedom (linear scale on the left and logarithmic on the right).
@@
[…]
@@ 5.2 Linear elasticity
As second problem we consider the linear elasticity equation
@@
on the domain $\Omega=(0,1)^{3}$ in the time interval $I=[0,2.5]$. By $\boldsymbol{u}$ we denote the displacement, $\boldsymbol{f}=(0,-1,0)^{T}$ is the right hand side vector and the stress tensor is given by
@@
with homogenous Dirichlet boundary conditions and zero initial conditions for $\boldsymbol{u}$ and $\boldsymbol{v}$.
We use this test case to illustrate the performance of the CUDA multigrid solver on adaptively refined meshes, see Fig. 5 for a visualization of the used mesh. Adaptive refinement is not driven in a problem specific way here but we consider typical cases of adaptive meshing, namely refinement towards a complete face of the box, refinement towards an edge, and refinement towards a vertex. This mimics resolving singularities that have a 2d pattern (face), a 1d pattern (edge) and a 0d pattern (vertex). Table 2 lists the number of mesh nodes and the fraction of mesh nodes that are hanging. The number of degrees of freedom is six times this number of mesh nodes as we have 3 deformation and 3 velocity components. Hanging nodes do not improve the approximation property but instead disrupt the structure of the problem by distorting the sparsity pattern of the matrix.
Figure 5: Visualization of the adaptive meshes. From left to right: Refinement towards one face, towards an edge, and towards one vertex. The fraction of hanging nodes and hence of unstructuredness increases from the left to the right.
@@
Table 2: Elasticity problem: Number of mesh nodes on different mesh levels (first sub-row) and fraction of nodes that are hanging nodes (second sub-row). From top to bottom we show the different adaptive refinement types, compare Fig. 5.
@@
We provide a detailed discussion of the intermediate case, the refinement towards one edge. Fig. 6(a) gives the timings on the finest mesh level (34 985 nodes, hence 209 910 degrees of freedom, about 25 600 of them in hanging nodes) for the CPU version and the GPU version. As for the transport-diffusion equation, times for initialization init and right hand side rhs do not change, as these parts are not implemented on the GPU. The computation time for the linear solver is, however, drastically reduced, compare Table 3(b). On the finest mesh, the solution time reduces from $1529\,\mathrm{s}$ to $96\,\mathrm{s}$, i.e. by a factor of 16. As the time for matrix assembly (init) and computing the right hand side (rhs) are negligible, the overall speedup is still close to 14 (compared to about 13 for the transport-diffusion problem on a much finer mesh).
Fig. 6(b) shows the scaling of the implementation with respect to the refinement. Note that the meshes are refined locally. Therefore, the number of multigrid levels increases in each step but the overall number of unknowns grows only slowly (not by a factor of 8 that would be expected on a 3d hexahedral mesh). Hence more and more unstructured features and hanging nodes appear on higher mesh levels. Due to the slow growth in the number of degrees of freedoms, the run-times on CPU and GPU increase only slowly with each refinement level, compare also Table 2. The results show very good efficiency of the GPU multigrid solver on locally refined meshes with no negative impact of having more and more hanging nodes and less regular structure in the problem.
@@
Figure 7 shows that the efficiency of the GPU parallelization is robust. On the finest mesh level belonging to the vertex-case in Table 3(c), the computational time of the solver only drops from $139\,\mathrm{s}$ on the CPU to $30\,\mathrm{s}$ on the GPU, a factor of about 4.5. However, this problem has just 25 000 unknowns and is too small to properly utilize a GPU. It is important to note that the GPU parallelization does not lead to an overhead at any point, which negatively affects the entire computing time.
@@ 5.3 Navier-Stokes equations
The error caused by mass lumping in the zeroth-order term (12) and the convective term (13) are of the same order as the polynomial approximation error and therefore do not effect the error asymptotics. This approach is sometimes denoted as ‘fully practical finite element method’ [3] as it allows for error estimates that reflect the full error including numerical quadrature.
@@ 5.3.1 Explicit projection solver
The fully discrete solution algorithm we employ solves the momentum equation explicitly and updates the pressure field by solving a Poisson problem (see Alg. 4). Hence, we introduce the solution space $S=H^{1}(\Omega)$ and its discrete counterpart $S_{h}$ that consists of $\mathbb{Q}_{1}$ polynomials on each cell $T$ of $\Omega_{h}$. These type of predictor-corrector methods for approximating incompressible flows requires an initial pressure field which must be calculated from the Poisson equation if the right hand side and initial velocity are not zero
@@
For justification and well-posedness of this equation, see [18]. Moreover, note that the velocity solutions $\boldsymbol{u}_{h}^{m}$ are so-called predictor solutions. The corrector velocities were eliminated as described in [16].
@@ 5.3.2 GPU realization
We also use custom kernels for efficiencly computing the products with the rectangular matrices corresponding to the discretization of the gradient of the pressure and the divergence of the velocity that is required in preprocessing.
@@ 5.3.3 Numerical Results
Refer to captionRefer to caption
Figure 8: Navier-Stokes problem: Wall-clock times CPU vs. GPU (GPU* is without custom kernels), logarithmic scale (left) and linear scale (right). Table 4 gives the raw numbers and it also depicts the time required for transfer of data between CPU and GPU.
@@
We split the timings into several components. momentum corresponds to Step 1 of Algorithm 4. Here, mom rhs is the assembly of the right hand side, which we further split into mom-rhs-nonlin, the multiplication with $C_{h}$, compare (15), into mom-rhs-p which is a multiplication with a rectangular matrix acting on the pressure and finally into mom-rhs-visc that covers the matrix-vector product representing the viscous term. Finally, mom-solve is the multiplication with the inverse mass matrix. The assembly of the right hand side is the most costly part in the algorithm. Porting it to the GPU reduces the computational cost by a factor of 10. Here, a substantial further reduction of computational time is reached by introducing custom kernels that preprocess the solution vector for assembling the outer products (14) and for handling rectangular matrices, compare Section 5.3.2. The reduction of computational time is due to saving memory transfer between CPU and GPU, compare columns GPU and GPU in Table 4.
pres covers Step 2 of Algorithm 4, the pressure Poisson problem. Again, we split the timings into pres-rhs for the assemble of the right hand side, a sparse matrix vector product computing the divergence of the velocity prediction and pres-solve which is the actual inversion of the Laplace problem using the GMRES multigrid solver.
pres-up finally corresponds to Step 3, the explicit update of the pressure. Here, pres-up-rhs is the assemble of the right hand side and pres-up-solve the inversion of the lumped pressure space mass matrix.
Memory is only transfered for initialization (first copy of matrices and vectors) as well as for the GMRES solver. Here however, only single floats, e.g. results of scalar products or short matrices used in the GMRES orthogonalization must be copied. Their dimension is in the order of number of GMRES steps (always less than 10).
@@ 6 Conclusions
In future work, we want to extend our GPU paralleization of Gascoigne 3d further. The most important step to reach the full flexibility is to realize the quadrature of variational formulations, i.e. the assembly of matrices and residuals, directly on the GPU. An interesting alternative to native CUDA is the use of the high-performance linear algebra libraries that are the backends of machine learning frameworks such as torch and jax, for example torch inductor or XLA. These provide flexible support for a range of accelerators, e.g. also TPUs, and also compilers that optimize the computations for the available compute hardware. Triton is also an interesting intermediate ground between native CUDA and higher level libraries.
The code of our implementation will be made available upon publication.
Acknowledgments
UK, ML and TR acknowledge the support of the GRK 2297 MathCoRe, funded by the Deutsche Forschungsgemeinschaft, Grant Number 314838170.

### v2

@@ 5.1 Transport-diffusion equation
Figure 4: Transport-diffusion equation: Wall-clock times CPU vs. GPU on the finest mesh level with 1 000 00010000001\,000\,0001 000 000 degrees of freedom (left: linear scale, right: log scale).
@@
We consider the backward Euler method with a time-step size $\Delta t=0.02$. Uniform quadrilateral spatial meshes with a range of sizes given by $h=2^{-n}$ where $n\in\{4,\dots,10\}$ are used to demonstrate the scaling efficiency of the GPU accelerated version. Hence, on the coarsest mesh the problem has $6414^{7}+2^{7+1}+1=16,$ 641 degrees of freedom whereas the finest mesh comprises $0004^{10}+2^{10+1}+1\approx 1,000,$ 000 degrees of freedom. In each of the $T/\Delta t=100$ time steps, a linear system must be solved. To ensure the robustness of the method, we use the GMRES solver, preconditioned with multigrid, to approximate the linear problems, cf. Sec. 4.1.5. Fig. 4 shows the wall-clock times comparing the CPU and GPU implementations on the finest mesh level. The labels init, rhs and solve refer to the initialization (mainly assembly of system matrix), to the computation of the right hand side, and to the actual solution, respectively. Matrix assembly and the computation of the right hand side involve numerical quadrature over the mesh elements which is not ported to the GPU. Hence, no speedup is observed, GPU timings can even be higher, as copying the matrix to the GPU involves some overhead, see also Fig. 5. The time for solving the linear problems, however, is reduced from $304.75\,\mathrm{s}$ on the CPU to $9.0\,\mathrm{s}$ on the GPU, a factor of about 34; see also Table 1 for the raw data. The geometric multigrid solver is robust on all meshes such that the number of linear steps varies only slightly between the mesh levels. For large problems, the non-accelerated assembly of the right hand side rhs hence becomes dominant. In Section 5.3 we describe how this term can be efficiently transferred to the GPU by using mass lumping.
@@ 5.2 Linear elasticity
Second, we consider the linear elasticity equation
@@
on the domain $\Omega=(0,1)^{3}$ in the time interval $I=[0,2.5]$. By $\boldsymbol{u}$ we denote the displacement, $\boldsymbol{f}=(0,-1,0)^{T}$ is the right hand side vector, the stress tensor is
@@
with homogenous Dirichlet boundary conditions and zero initial conditions for $\boldsymbol{u}$ and $\boldsymbol{v}$. We use this test case to illustrate the performance of the CUDA multigrid solver on adaptively refined meshes, see Fig. 6 for a visualization of the used mesh. Adaptive refinement is not driven in a problem specific way here but we consider typical cases of adaptive meshing, namely refinement towards a complete face of the box, refinement towards an edge, and refinement towards a vertex. This mimics resolving singularities that have a 2d pattern (face), a 1d pattern (edge) and a 0d pattern (vertex). Table 2 lists the number of mesh nodes and the fraction of mesh nodes that are hanging. The number of degrees of freedom is six times this number of mesh nodes as we have 3 deformation and 3 velocity components. Hanging nodes do not improve the approximation property but instead disrupt the structure of the problem by distorting the sparsity pattern of the matrix.
Figure 6: Visualization of the adaptive meshes. From left to right: Refinement towards one face, towards an edge, and towards one corner. The fraction of hanging nodes serving as measure for the unstructuredness increases from the left to the right.
@@
Table 2: Elasticity problem: Number of mesh nodes on different mesh levels (first sub-row) and fraction of nodes that are hanging nodes (second sub-row). From top to bottom we show the different adaptive refinement types, see Fig. 6.
@@
We provide a detailed discussion of the intermediate case, the refinement towards one edge. Fig. 7(a) gives the timings on the finest mesh level (34 985 nodes, hence 209 910 degrees of freedom, about 25 600 of them in hanging nodes) for the CPU version and the GPU version. As for the transport-diffusion equation, times for initialization init and right hand side rhs do not change, as these parts are not implemented on the GPU. The computation time for the linear solver is, however, drastically reduced, see Table 3(b). On the finest mesh, the solution time reduces from $306\,\mathrm{s}$ to $15\,\mathrm{s}$, i.e. by a factor of 20. As the time for matrix assembly (init) and computing the right hand side (rhs) are negligible, the overall speedup is still close to 17 (compared to about 8 for the transport-diffusion problem on a much finer mesh). As for the transport-diffusion problem, the initialization (init) may require more time on the GPU as this includes the cost for copying the matrix from the CPU to the GPU. Also, since the effort is negligible for this problem, the right hand side has been assembled on the CPU for simplicity. Therefore, the timinigs on the GPU include overhead for copying the data to the GPU.
Fig. 7(b) shows the scaling of the implementation with respect to the refinement. Note that the meshes are refined locally. Therefore, the number of multigrid levels increases in each step but the overall number of unknowns grows only slowly (not by a factor of 8 that would be expected on a 3d hexahedral mesh). Hence more and more unstructured features and hanging nodes appear on higher mesh levels. Due to the slow growth in the number of degrees of freedoms, the run-times on CPU and GPU increase only slowly with each refinement level, see also Table 2. The results show very good efficiency of the GPU multigrid solver on locally refined meshes with no negative impact of having more and more hanging nodes and less regular structure in the problem.
@@
Figure 8 shows that the efficiency of the GPU parallelization is robust. On the finest mesh level belonging to the vertex-case in Table 3(c), the computational time of the solver only drops from $29\,\mathrm{s}$ on the CPU to $8\,\mathrm{s}$ on the GPU, a factor of about 3.5. However, this problem has just 25 000 unknowns and is too small to properly utilize a GPU. It is important to note that the GPU parallelization does not lead to an overhead at any point, which negatively affects the entire computing time.
@@ 5.3 Navier-Stokes equations
The error caused by mass lumping in the zeroth-order term (12) and the convective term (13) are of the same order as the polynomial approximation error and therefore do not affect the error asymptotics. This approach is sometimes denoted as ‘fully practical finite element method’ [3] as it allows for error estimates that reflect the full error including numerical quadrature. We have recently published the complete error analysis for an explicit pressure correction method based on this approximation [22]. Both theoretical analysis and numerical examples show that the explicit treatment based on the interpolation of the nonlinearity does not lead to any additional error.
@@ 5.3.1 Explicit projection solver
The fully discrete solution algorithm we employ solves the momentum equation explicitly and updates the pressure field by solving a Poisson problem (see Alg. 4). Hence, we introduce the solution space $S=H^{1}(\Omega)$ and its discrete counterpart $S_{h}$ that consists of $\mathbb{Q}_{1}$ polynomials on each cell $T$ of $\Omega_{h}$. This type of predictor-corrector methods for approximating incompressible flows requires an initial pressure field which must be calculated from the Poisson equation if the right hand side and initial velocity are not zero
@@
For justification and well-posedness of this equation, see [20]. Moreover, note that the velocity solutions $\boldsymbol{u}_{h}^{m}$ are so-called predictor solutions. The corrector velocities were eliminated as described in [18]. The complete error analysis for this fully explicit variant of a pressure correction scheme is found in [22] and its extension to the Boussinesq approximation is described in [13].
@@ 5.3.2 GPU realization
We also use custom kernels to efficiently compute the products with the rectangular matrices corresponding to the discretization of the gradient of the pressure and the divergence of the velocity that is required in preprocessing. Appendix A gives details on the implementation and in particular on the required changes in the user code to employ GPU parallelization and on the integration of custom kernels.
@@ 5.3.3 Numerical Results
Figure 9: Navier-Stokes problem: Wall-clock times CPU vs. GPU (GPU* is without custom kernels), log scale (left) and linear scale (right). Table 4 gives the raw numbers and it also depicts the time required for transfer of data between CPU and GPU.
@@
We split the timings into several components. momentum corresponds to Step 1 of Algorithm 4. Here, mom-rhs is the assembly of the right hand side, which we further split into mom-rhs-nonlin, the multiplication with $C_{h}$, see (15), into mom-rhs-p which is a multiplication with a rectangular matrix acting on the pressure and finally into mom-rhs-visc that covers the matrix-vector product representing the viscous term. By mom-solve we denote the multiplication with the inverse mass matrix. The assembly of the right hand side combined in mom-rhs is the most costly part in the solution of the momentum equation. Porting it to the GPU reduces the computational cost by a factor of 6.
The column denoted as GPU* is purely based on cuSparse and no custom kernels are used. In particular, cuSparse does not provide functionality to compute the node-wise products (14) which are therefore still computed on the CPU. This causes the substantial costs for the memory transfer listed as copy. By introducing a custom CUDA kernel to directly compute (14) on the GPU, a substantial further reduction of computational time is achieved, see also Section 5.3.2. A further custom kernels is used to optimize the transfer between the scalar pressure space and the vector-valued velocity space required in mom-rhs-p. The column labelled GPU lists the timings based on these additional custom kernels. The costs for memory transfer and hence the overall runtime is hereby substantially reduced giving a speedup (for the momentum part momentum) by a factor of nearly 40, see Table 4.
pres covers Step 2 of Algorithm 4, the pressure Poisson problem. Again, we split the timings into pres-rhs for the assemble of the right hand side, a sparse matrix vector product computing the divergence of the velocity prediction and pres-solve which is the actual inversion of the Laplace problem using the GMRES multigrid solver. pres-up finally corresponds to Step 3, the explicit update of the pressure. Here, pres-up-rhs is the assemble of the right hand side and pres-up-solve the inversion of the lumped pressure space mass matrix.
Both pres-rhs and pres-up.rhs require matrix-vector multiplications with rectangular matrices. The timings listed under GPU* require a memory transfer to the CPU to map between pressure and velocity space (similar to mom-rhs-p described above). Acceleration is achieved by again using a custom kernel to directly work on the GPU.
@@
5.4 Limitations
Figure 10: GPU memory usage depending on the degrees of freedom of the solution. The problem size is increased by repeated mesh refinement as described in Sections 5.1 and 5.2. For Navier-Stokes, the mesh is also refined uniformly. Dashed lines are linear extrapolations from the respective largest three sizes. Dotted lines indicate the available memory on 3 generations of data center GPUs.
In order to achieve a significant speedup with the GPU, all the required data needs to be kept on the device. Because of the smaller size of GPU-RAM compared to CPU-RAM, the maximum problem size that can be solved with the GPU implementation is more limited. The device memory usage of the presented examples is shown in Fig. 10. On an H100 GPU, the limit for the 3D Navier-Stokes problem is of the order $10^{7}$. One potential direction to address this is a multi-GPU system. A challenge is thereby to load-balance between the GPUs with non-uniform mesh refinement. We plan to investigate this in more detail in future work.
For speed, the bottleneck of the GPU implementation is the memory bandwidth. The increase in bandwidth by a factor of 10 from the DDR4 RAM accessed by the CPU to the HBMe2 memory utilized on the H100 GPU (2 TB/s) is complemented by a similar increase in potential compute throughput. In the Navier-Stokes example, roughly $2/3$ of the GPU cycles are spend on sparse matrix-vector products, which have an arithmetic intensity of 0.39FLOP/byte (0.15FLOP/byte for the transposed variant) and therefore achieve just $\frac{1}{32}$ ($\frac{1}{83}$) of the peak double-precision performance. As part of cuSPARSE, the kernels are already highly optimized with a memory throughput of $91\%$ ($82\%$) and the low arithmetic intensity is inherent to the computational problem. Major improvements would therefore require fusing some of the roughly 3500 kernel calls needed for a single time-step together in order to reuse the data. This is not practical with the current approach that heavily leverages the cuSPARSE library. All the custom kernels together use less than $3\%$ of the GPU time. While some of these could be eliminated by assembling rectangular block matrices in Gascoigne 3D, the speedup would therefore be negligible. The most promising direction for further optimizations is therefore a switch to single precision, since this effectively doubles the arithmetic intensity.
@@ 6 Conclusions
We are planning several steps for further optimisation: A major limitation of the previous GPU parallelization is that the matrices are assembled on the CPU and then transferred statically to the GPU. This severely restricts the applicability to nonlinear problems. The obvious option is to assemble the matrices directly on the GPU. However, this is complicated by a large number of options in the choice and control of the discretization and the use of adaptive grids. Alternatively, we plan to use the GPU and CPU in hybrid mode, so that for nonlinear problems the Jacobi matrix of the Newton solver is always built in the background on the CPU and then transferred to the GPU. This can be seamlessly integrated in the usual inexact Newton algorithm, which only reassembles the Jacobian when the convergence rate deteriorates. A further important step is the GPU parallelization of more powerful multigrid smoothers than the block Jacobi iteration. The realisation of a Vanka smoother can be easily implemented with custom kernels and then allows the direct solution of saddle point problems.
An interesting alternative to native CUDA is the use of the high-performance linear algebra libraries that are the backends of machine learning frameworks such as torch and jax, for example torch inductor or XLA. These provide flexible support for a range of accelerators, e.g. also TPUs, and also compilers that optimize the computations for the available compute hardware. Triton is also an interesting intermediate ground between native CUDA and higher level libraries.
Funding
UK, ML and TR acknowledge the support of the GRK 2297 MathCoRe, funded by the Deutsche Forschungsgemeinschaft, Grant Number 314838170. RJ acknowledges that support was provided by Schmidt Sciences.
Code availability
The source code of the Gascoigne library with CUDA support is found in the Zenodo repository https://zenodo.org/records/13891228 [26]. This repository includes all scripts required to reproduce the examples discussed in this manuscript.
Appendix A Details on the implementation
To describe the necessary changes to the source code for using GPU acceleration and also to explain the role of the custom kernels we give details on the implementation of the interpolated nonlinearity in (15). We start with the basic CPU Gascoigne version:
Assembling the nonlinearity (CPU)
void nonlinear(Vector& nl, const Vector& u) {
#pragma omp parallel for
for (size_t i = 0; i < n; ++i) {
nl(i,0) = u(i,0)*u(i,0); nl(i,1) = u(i,0)*u(i,1);
nl(i,2) = u(i,0)*u(i,2); nl(i,3) = u(i,1)*u(i,1);
nl(i,4) = u(i,1)*u(i,2); nl(i,5) = u(i,2)*u(i,2); }
void Nonlinear(Vector& f, const Vector& u) {
nonlin(nl, u); // assembles node-wise product
GetSolver()->vmulteq(C, f, nl, -1.); // f = -C*nl
The function nonlinear(...) has no native implementation in CuSparse. Hence, in the GPU* version, where only sparse matrix vector operations are run on the GPU, the following minimal modification of the function Nonlinear(...) is needed:
Assembling the nonlinearity (GPU*)
void Nonlinear(Vector& f, const Vector& u) {
DeactivateCuda(u); // copies vectors to CPU
nonlin(nl, u); // assembles node-wise product
ActivateCuda(nl); // copies result to GPU
GetSolver()->vmulteq(C, f, nl, -1.); // f = -C*nl
The two functions ActivateCuda(...) and DeactivateCuda(...) set internal flags to activate or deactivate matrix-vector handling on the GPU and they transfer the data of the corresponding vectors such that the function nonlin(...) can be processed on the CPU. To avoid this data transfer we must shift this function to a custom kernel.
Assembling the nonlinearity (GPU)
__global__ void
nonlin_device(size_t n, double* nl, double* u) {
if (threadIdx.x + blockIdx.x * blockDim.x >= n)
return;
nl[id*6+0]=u[id*3+0]*u[id*3+0];
nl[id*6+1]=u[id*3+0]*u[id*3+1];
nl[id*6+2]=u[id*3+0]*u[id*3+2];
nl[id*6+3]=u[id*3+1]*u[id*3+1];
nl[id*6+4]=u[id*3+1]*u[id*3+2];
nl[id*6+5]=u[id*3+2]*u[id*3+2];
void nonlin(IndexType n, double* dest, double* src) {
size_t tpb = 1024;
size_t blocks = max(1, ceil(n / tpb))
nonlin_device$<$blocks, tpb$>$(n, dest, src);
void Nonlinear(Vector& f, const Vector& u) {
nonlin(u.n(), nl, u); // node-wise product on GPU
GetSolver()->vmulteq(C, f, nl, -1.); // f = -C*nl